OpenStudy (anonymous):

One number is 2 more than another. The difference between their squares is 24. What are the numbers?

OpenStudy (freckles):

say the numbers they are talking about is x and y. x is 2 more than y <--translate this into math symbols

OpenStudy (anonymous):

so two equtions? x+y=24 2x+2y=24

OpenStudy (freckles):

well you need to translate the following two sentences into mathematical equaitons: "x is 2 more than y" "the difference between the square of x and the square of y is 24.

OpenStudy (freckles):

x is 2 more than y is means equals more than means +

OpenStudy (freckles):

 x=2+y